If the equation for the demand curve is Q=45-2p and the equation for the supply curve is Q=-21+4p then the equilibrium values for price and quantity are?

1
Expert's answer
2022-03-24T16:13:53-0400

"Q"d "=45-2p"

"Q"s "= -21+4p"


At equilibrium "Q"d "=Q"s

"45-2p=-21+4p"

"45+21=4p+2p"

"66= 6p"

"\frac{66}{6}" "=\frac{6p}{6}"

"\therefore p=11"

"Q"d "=45-2(11)"

Qd "=" "23"

At equilibrium the price is 11 and the equilibrium quantity is 23.